# SKELETON-BASED SIGN LANGUAGE RECOGNITION USING A DUAL-STREAM SPATIO-TEMPORAL DYNAMIC GRAPH CONVOLUTIONAL NETWORK

### v1

@@ Abstract
Isolated Sign Language Recognition (ISLR) is challenged by gestures that are morphologically similar yet semantically distinct, a problem rooted in the complex interplay between hand shape and motion trajectory. Existing methods, often relying on a single reference frame, struggle to resolve this geometric ambiguity. This paper introduces Dual-SignLanguageNet (DSLNet), a dual-reference, dual-stream architecture that decouples and models gesture morphology and trajectory in separate, complementary coordinate systems. Our approach utilizes a wrist-centric frame for view-invariant shape analysis and a facial-centric frame for context-aware trajectory modeling. These streams are processed by specialized networks—a topology-aware graph convolution for shape and a Finsler geometry-based encoder for trajectory—and are integrated via a geometry-driven optimal transport fusion mechanism. DSLNet sets a new state-of-the-art, achieving 93.70%, 89.97% and 99.79% accuracy on the challenging WLASL-100, WLASL-300 and LSA64 datasets, respectively, with significantly fewer parameters than competing models.
@@ 2.1 Dual-Reference Frame Normalization
We decouple the input into two complementary representations by normalizing the skeleton sequence with respect to two different reference frames.
Wrist Morphological Frame. To capture the intrinsic shape of the hand, we normalize the hand joints $\mathcal{J}$ relative to the wrist joint $h_{w}$. This creates a view-invariant representation of the hand’s morphology.
$\mathcal{X}^{\prime}_{\text{shape}}(t,i)=h_{i}(t)-h_{w}(t),\quad i\in\mathcal{J}$ (1)
Facial Semantic Frame. To model the hand’s trajectory relative to the body, we normalize the wrist’s position with respect to a stable facial reference frame. This frame is defined by the centroid $c_{f}(t)$ and scale $s_{f}(t)$ of key facial landmarks $\mathcal{K}_{f}$. This captures the crucial spatial context of the gesture.
$\mathcal{X}^{\prime}_{\text{traj}}(t)=\frac{h_{w}(t)-c_{f}(t)}{s_{f}(t)+\varepsilon}$ (2)
@@ 3.1 Implementation Details
Datasets. We evaluate DSLNet on three widely-used datasets: LSA64, WLASL-100, and WLASL-300.
Preprocessing. For each video, we extract 21 hand keypoints and 5 facial keypoints (nose and four mouth corners) using MediaPipe. The skeleton coordinates are normalized to the range [1,1][-1,1]. We apply standard data augmentation, including random rotation, scaling, Gaussian noise, and temporal stretching to prevent overfitting.
Training. The model is implemented in PyTorch and trained on a single NVIDIA RTX 4090 GPU. We use the AdamW optimizer with a cosine annealing learning rate schedule.
@@ 3.3 Ablation Studies
Effect of Dual-Reference Frames. Table 2 validates the necessity of our dual-reference strategy. While using a single reference frame like ”Wrist Morphological Only” provides a strong baseline (e.g., 92.71% on WLASL-100), it remains suboptimal. By combining it with the ”Facial Semantic Frame”, our full model achieves significant accuracy gains across all datasets. Specifically, on WLASL-100 and WLASL-300, the dual-frame approach improves accuracy by 0.99% and 0.46% respectively, compared to the best way of others. This confirms that the two frames offer complementary geometric information essential for resolving ambiguities.
@@
Global Normalization Only 99.32 93.05 89.43
Wrist Morphological Only 99.15 92.71 89.51
Facial Semantic Only 99.47 92.56 88.34
@@
Dual-stream + Concatenation 93.45 90.01 86.56
@@ 4 Conclusion
This paper introduced DSLNet, a framework that resolves geometric ambiguities in skeleton-based sign language recognition by decoupling morphology and trajectory analysis. The synergy between a topology-aware shape network, a direction-sensitive trajectory encoder, and a geometry-driven optimal transport fusion mechanism leads to state-of-the-art performance. DSLNet achieves superior accuracy on challenging benchmarks with remarkable parameter and computational efficiency. This work validates the importance of multi-reference geometric modeling, offering a robust and practical solution for real-world ISLR. Future work will extend this framework to continuous sign language recognition.

### v2

@@ Abstract
Isolated Sign Language Recognition (ISLR) is challenged by gestures that are morphologically similar yet semantically distinct, a problem rooted in the complex interplay between hand shape and motion trajectory. Existing methods, often relying on a single reference frame, struggle to resolve this geometric ambiguity. This paper introduces Dual-SignLanguageNet (DSLNet), a dual-reference, dual-stream architecture that decouples and models gesture morphology and trajectory in separate, complementary coordinate systems. The architecture processes these streams through specialized networks: a topology-aware graph convolution models the view-invariant shape from a wrist-centric frame, while a Finsler geometry-based encoder captures the context-aware trajectory from a facial-centric frame. These features are then integrated via a geometry-driven optimal transport fusion mechanism. DSLNet sets a new state-of-the-art, achieving 93.70%, 89.97%, and 99.79% accuracy on the challenging WLASL-100, WLASL-300, and LSA64 datasets, respectively, with significantly fewer parameters than competing models.
@@ 2.1 Dual-Reference Frame Normalization
To resolve geometric ambiguities, we decouple the raw skeleton sequence into two semantically distinct representations. This is achieved by projecting the keypoint data onto two complementary, dynamically defined coordinate systems.
Wrist Morphological Frame. To isolate the intrinsic hand configuration from global motion, we apply a translation operator $\mathcal{T}$ that re-centers the hand joint coordinates into a local, wrist-centric frame. This operator, parameterized by the wrist’s position $h_{w}(t)$, generates a translation-invariant representation $\mathcal{X}^{\prime}_{\text{shape}}$ that models pure morphology:
$\mathcal{X}^{\prime}_{\text{shape}}(t,i)=\mathcal{T}_{h_{w}(t)}(h_{i}(t)),\quad i\in\mathcal{J}$ (1)
where $\mathcal{T}_{\mathbf{p}}(\mathbf{x})=\mathbf{x}-\mathbf{p}$ defines the translation.
Facial Semantic Frame. To capture the gesture’s contextual trajectory, we employ a normalization operator $\mathcal{N}$ that projects the wrist’s global motion into a dynamic, body-centric coordinate system. This operator is anchored by a time-varying origin $c_{f}(t)$ and scale factor $s_{f}(t)$ derived from facial landmarks. The resulting representation, $\mathcal{X}^{\prime}_{\text{traj}}$, preserves crucial spatial semantics while achieving pose and scale invariance:
$\mathcal{X}^{\prime}_{\text{traj}}(t)=\mathcal{N}_{c_{f}(t),s_{f}(t)}(h_{w}(t))$ (2)
where $\mathcal{N}_{\mathbf{c},s}(\mathbf{x})=(\mathbf{x}-\mathbf{c})/(s+\varepsilon)$ defines the normalization transformation.
@@ 3.1 Implementation Details
Datasets. We evaluate DSLNet on three widely-used datasets: LSA64[10] and WLASL[11](WLASL-100 and WLASL-300).
Preprocessing. For each video, we extract 21 hand keypoints and 5 facial keypoints (nose and four mouth corners) accurately. The skeleton coordinates are normalized to the range [1,1][-1,1]. We apply standard data augmentation, including random rotation, scaling, Gaussian noise, and temporal stretching to prevent overfitting.
Training. The model is implemented in PyTorch and trained on the NVIDIA RTX 4090 GPU. We use the AdamW optimizer with a cosine annealing learning rate schedule.
@@ 3.3 Ablation Studies
Effect of Dual-Reference Frames. Table 2 validates our dual-reference strategy against a global normalization baseline. Our approach consistently yields superior performance, securing gains of 0.38% on WLASL-100 and 0.74% on WLASL-300. This confirms that decomposing gestures into view-invariant shape and context-aware trajectory provides a more potent feature space, effectively resolving geometric ambiguities for a more robust representation.
@@
Global Normalization 99.15 93.32 89.23
@@
Dual-stream + Concatenation 92.45 90.01 86.56
@@ 4 Conclusion
This paper introduced DSLNet, a framework that resolves geometric ambiguities in skeleton-based sign language recognition by decoupling morphology and trajectory analysis. The synergy between a topology-aware shape network, a direction-sensitive trajectory encoder, and a geometry-driven optimal transport fusion mechanism leads to state-of-the-art performance. DSLNet achieves superior accuracy on challenging benchmarks with remarkable parameter and computational efficiency. This work validates the importance of multi-reference geometric modeling, offering a robust and practical solution for real-world ISLR. Future work will extend this framework to continuous sign language recognition and reduce dependency on precise keypoint detection.